### v1

Abstract
Maximal cliques play a fundamental role in numerous application domains, where their enumeration can prove extremely useful. Yet their sheer number, even in sparse real-world graphs, can make them impractical to be exploited effectively. To address this issue, one approach is to enumerate $\ell$-isolated maximal cliques, whose vertices have (on average) less than $\ell$ edges toward the rest of the graph. By tuning parameter $\ell$, the degree of isolation can be controlled, and cliques that are overly connected to the outside are filtered out. Building on Tomita et al.’s very practical recursive algorithm for maximal clique enumeration, we propose four pruning heuristics, applicable individually or in combination, that discard recursive search branches that are guaranteed not to yield $\ell$-isolated maximal cliques. Besides proving correctness, we characterize both the pruning power and the computational cost of these heuristics, and we conduct an extensive experimental study comparing our methods with Tomita’s baseline and with a state-of-the-art approach. Results show that two of our heuristics offer substantial efficiency improvements, especially on real-world graphs with social network properties.
keywords:
Maximal cliques, isolated cliques, enumeration algorithms, clique summarization, algorithm engineering
category:
[…]
1 Introduction
Cliques capture dense, highly-connected network structures. Their enumeration is important in many application domains, such as Web mining, bioinformatics, computational chemistry, social and financial network analysis. In particular, enumerating maximal cliques — those that cannot be extended by another vertex — helps reveal the graph’s structure by uncovering tightly connected groups of vertices that share meaningful relationships.
From both theoretical and practical perspectives, the problem presents significant challenges. The set of maximal cliques includes indeed also maximum cliques, but computing a maximum clique is a well-known $\mathsf{NP}$-hard problem [DBLP:conf/coco/Karp72], as well as $\mathsf{W}$[1]-hard [DBLP:series/mcs/DowneyF99, downey2012parameterized], and hard to approximate within a factor of $n^{1-\epsilon}$ [hastad1999clique, DBLP:conf/stoc/Zuckerman06]. In general, the number of maximal cliques in an $n$-vertex graph can be exponential in $n$ [Moon1965OnCI]. Hence, not only their enumeration requires considerable time, but navigating such a large number of structures is also difficult. Imposing additional constraints may pinpoint a subset of domain-relevant structures, substantially reduce the number of enumerated cliques, and enable more effective utilization. We refer to [DFP25] for a taxonomic survey of different summarization approaches.
In contexts where cliques with weak connections to the rest of the graph are more meaningful than those strongly intertwined with it, the concept of $\ell$-isolation may prove particularly useful. First introduced in [DBLP:conf/esa/ItoIO05, DBLP:journals/talg/ItoI09], $\ell$-isolation requires that the number of edges leaving a clique of size $k$ does not exceed $\ell\cdot k$. In the same works, the authors proposed an $\mathsf{FPT}$ algorithm in the isolation factor $\ell$ to enumerate all $\ell$-isolated maximal cliques. It was also shown that $\ell$ must be constant to achieve linear-time enumeration, since otherwise one can construct instances admitting a superlinear number of $\ell$-isolated maximal cliques.
In [DBLP:journals/tcs/KomusiewiczHMN09], two additional isolation principles were introduced, namely max- and min-isolation. A systematic comparison of different isolation notions, complemented by an experimental evaluation on synthetic and financial networks, was presented in [DBLP:journals/tcs/HuffnerKMN09], underscoring the importance of isolated clique enumeration for network analysis. Isolation concepts have been also explored in temporal networks [DBLP:journals/netsci/MolterNR21, DBLP:conf/gpc/GaoHYYM20], and applied to practical scenarios such as web structure mining [DBLP:journals/ieicet/UnoOU07], hierarchical representations of scale-free networks [DBLP:journals/jgaa/ShigezumiUW11], and network contraction via isolated cliques [DBLP:journals/ieicet/UnoO13]. Alternative isolation models have also been proposed, including $\tau$-isolation based on coreness [DBLP:conf/mldm/OkuboHT16], as well as variants for stars [DBLP:conf/waw/UnoOU06], bicliques [DBLP:journals/ijcm/AlamgirKH17], and even approaches inspired by quantum algorithms [DBLP:conf/opodis/GallNNO24].
Our contribution. In this paper, building upon the well-known Bron-Kerbosch algorithm [DBLP:journals/cacm/BronK73] for maximal clique enumeration, improved by a pivoting strategy introduced by Tomita et al. [DBLP:conf/cocoon/TomitaTT04, DBLP:journals/tcs/TomitaTT06], we first design four pruning heuristics aimed at reducing the search space, proving their correctness. The heuristics, that can be applied individually or in combination, allow the enumeration algorithm to focus only on promising parts of the search tree, discarding branches that are guaranteed not to yield $\ell$-isolated maximal cliques. Following a theoretical characterization of the heuristics’ pruning power and computational cost, which turn out to be inversely proportional, we conduct an extensive experimental study assessing their practical impact. Our analysis addresses both real and synthetic instances and contributes along several dimensions:
  • It sheds light on the properties of $\ell$-isolated cliques relative to the full set of maximal cliques in real-world graphs. This can be particularly useful for guiding the choice of parameter $\ell$.
  • It gives insights into the relative power of our pruning strategies and of their combinations, identifying the most effective ones.
  • It compares our approach with a carefully tuned baseline algorithm [DBLP:conf/cocoon/TomitaTT04, DBLP:journals/tcs/TomitaTT06] and with a top-down state-of-the-art approach to the problem [DBLP:journals/tcs/HuffnerKMN09]. Overall, two of our variants offer substantial efficiency improvements over the state of the art, by as much as a factor of 4, especially when used on real-world graphs that exhibit social network properties.
Paper organization. The rest of this paper is organized as follows. We provide preliminary definitions and a description of the baseline enumeration algorithm in Section˜2. The pruning strategies are introduced and theoretically analyzed in Section˜3. Our experimental setup and the main outcomes of our analysis are discussed in Section˜4 and Section˜5, respectively. We summarize our results and outline open problems in Section˜6.
2 Preliminaries
In this section we first introduce preliminary definitions and properties of $\ell$-isolation (Section˜2.1), and then describe the maximal clique enumeration algorithm underlying our approach (Section˜2.2). Throughout the paper we consider finite simple graphs. For a graph $G$, let $V=V(G)$ and $E=E(G)$ denote its vertex and edge sets, with $n=|V|$ and $m=|E|$ their respective sizes. For a vertex $v$, let $\delta(v)$ denote its degree and $N(v)=\{u\mid(u,v)\in E\}$ its set of neighbors. Clearly, $\delta(v)=|N(v)|$. For any vertex subset $I\subseteq V$, let $G[I]$ denote the subgraph of $G$ induced by $I$, with vertex set $V[I]=I$ and edge set $E[I]=\{(u,v)\in E\mid u,v\in I\}$ consisting of edges with both endpoints in $I$. For a vertex $v\in I$, let $\delta_{e}(v,I)$ denote its external degree with respect to $I$, i.e., $\delta_{e}(v,I)=|\{(u,v)\in E\mid u\notin I\}|$. The external degree of $I$, denoted as $\delta_{e}(I)$, is then the number of edges with one endpoint in $I$ and the other in $V\setminus I$, i.e.,
$\delta_{e}(I)=\sum_{v\in I}\delta_{e}(v,I)=|\{(u,v)\in E\mid v\in I\penalty 10000\ \mbox{and}\penalty 10000\ u\in V\setminus I\}|$
Moreover, given $I,L\subseteq V$ with $I\cap L=\varnothing$, the external degree of $I$ neglecting $L$, denoted as $\delta_{e}(I,L)$, is the number of edges with one endpoint in $I$ and the other in $V\setminus(I\cup L)$.
2.1 $\ell$-isolated cliques
A clique is a complete subgraph of $G$, and is maximal if it is not contained in any larger clique. The concept of $\ell$-isolation has been first introduced in [DBLP:journals/talg/ItoI09]: for any integer value $\ell>0$, a set $I\subseteq V$ of vertices $G$ of size $|I|=k$ is $\ell$-isolated in $G$ if it has less than $\ell\cdot k$ outgoing edges, i.e., $\delta_{e}(I)<\ell\cdot k$. In particular, we are interested into maximal cliques that are also $\ell$-isolated ($\ell$-isolated maximal cliques). We notice that $I$ may contain vertices with more than $\ell$ external edges: the property of $\ell$-isolation holds on average over the vertices of $I$, implying that the average external degree of vertices in $I$ is less than $\ell$ [^1] A stronger notion of max-$\ell$-isolation requires every $v\in I$ to have fewer than $\ell$ external edges, but will not be considered in this paper. Algorithmic techniques similar to those presented here could nevertheless be applied..
We remark that the property of $\ell$-isolation is monotone. Namely, if $\ell_{1}>\ell_{2}>0$, then any $\ell_{2}$-isolated set is also $\ell_{1}$-isolated. This follows immediately from the definition of $\ell$-isolation, since for an $\ell_{2}$-isolated set $I$ we have $\delta_{e}(I)<\ell_{2}\cdot k<\ell_{1}\cdot k$ when $\ell_{1}>\ell_{2}$. Hence, when searching for $\ell$-isolated maximal cliques, the larger is the value of $\ell$, the weaker is the isolation of the cliques relative to the rest of the graph, and the is larger the number of $\ell$-isolated cliques in the graph. If $\ell$ is high enough, the set of $\ell$-isolated maximal cliques coincides with the set of all maximal cliques of $G$.
2.2 Maximal clique enumeration with pivoting
Our approach hinges upon the recursive backtracking procedure for maximal clique enumeration outlined in Algorithm˜1. The procedure extends the well-known algorithm by Bron and Kerbosch [DBLP:journals/cacm/BronK73] with a pivoting strategy introduced in [DBLP:conf/cocoon/TomitaTT04, DBLP:journals/tcs/TomitaTT06], aimed at reducing the time spent on non-maximal cliques. Three sets of vertices are maintained by procedure ProcMCE: $C$ represents the clique under construction, $P$ contains candidate vertices that can potentially extend $C$, and $X$ holds vertices that should no longer be considered, as they are already part of maximal cliques obtained from $C$ or its subsets. As an invariant property, vertices in $P$ and $X$ are adjacent to all vertices in $C$. The idea is to add at each step a vertex, selected from the candidates in $P$, expanding the current clique $C$ until it becomes maximal. At the first invocation, $C=X=\varnothing$ and $P=V$. In procedure ProcMCE, base cases occur when $P$ is empty: then $C$ cannot be further expanded and is returned only if $X=\varnothing$, ensuring its maximality. Otherwise, a pivot vertex $v_{p}$ is chosen so as to minimize recursive calls: ProcMCE is called on each $v\in P$ not adjacent to $v_{p}$ (possibly including $v_{p}$). Correctness follows since, for each neighbor $w$ of $v_{p}$, any maximal clique containing $C\cup\{w\}$ will still be reported by recursing on either $C\cup\{v_{p}\}$ or $C\cup\{u\}$, where $u\in P$ is a non-neighbor of $v_{p}$. Recursive calls update $(C,P,X)$ to $(C\cup\{v\},\,P\cap N(v),\,X\cap N(v))$, ensuring vertices in $P\cup X$ to remain adjacent to $C$. Afterwards, $v$ is removed from further consideration.
Algorithm 1 TomitaTanakaTakahashi
1:Input: Graph $G=(V,E)$
2:Output: The set $\mathcal{M}(G)$ of all maximal cliques of $G$
[…]
4:
5:procedure ProcMCE($C,P,X$)
6:  if $P=\varnothing$ and $X=\varnothing$ then return $C$
7:  Choose a pivot vertex $v_{p}$ in $P\cup X$ with highest $|N(v_{p})\cap P|$
8:  for all $v\in P\setminus N(v_{p})$ do
[…]
10:   $P\leftarrow P\setminus\{v\}$
11:   $X\leftarrow X\cup\{v\}$
When invoked on a graph $G$, Algorithm˜1 implicitly explores a search tree $\mathcal{T}_{G}$. Each node of $\mathcal{T}_{G}$ corresponds to a recursive call and is uniquely associated with a triple $(C,P,X)$. The leaves of $\mathcal{T}_{G}$ are nodes of the form $(C,\varnothing,X)$. If $X=\varnothing$, the leaf corresponds to a maximal clique that is enumerated. Otherwise, the maximality condition is not satisfied (vertices in $X$ could be added to $C$), the clique is discarded, but no recursive call needs to be done.
Algorithm˜1 requires $\mathcal{O}(3^{n/3})$ time [DBLP:journals/tcs/TomitaTT06]. While being optimal in the worst-case (matching the number of maximal cliques in Moon-Moser graphs [Moon1965OnCI]), it has $\Omega(3^{n/6})$ delay and no polynomial-delay variant can exist, even with different pivoting, unless ${\mathsf{P}}={\mathsf{NP}}$ [DBLP:journals/tcs/ConteT22]. Despite these worst-case bounds, Bron-Kerbosch-based approaches are known to perform very efficiently in practice [DBLP:journals/jea/EppsteinLS13, DBLP:journals/tcs/TomitaTT06].
3 Pruning for $\ell$-isolated maximal clique enumeration
In this section we show how to adapt Algorithm˜1 to enumerate only $\ell$-isolated maximal cliques, introducing different pruning strategies to cut uninteresting branches of the search tree. We also analyze the pruning capabilities and computational cost of each strategy, showing an inverse relation that suggests a natural order in which the heuristics may be applied in combination.
Growing $\ell$-isolated cliques. We recall from Section˜2.2 that in Algorithm˜1, during the enumeration, $C$ and $P$ are the current candidate clique and the perspective set, respectively. Let $T$ be any maximal clique in the subgraph $G[P]$ induced by $P$. Thanks to the invariant property that $P$ is always fully connected to $C$ during the execution, it is not difficult to see that $C\cup T$ is a clique in $G$, though not necessarily $\ell$-isolated. Our approach, when processing a node $r$ of the search tree $\mathcal{T}_{G}$ associated with a triple $(C,P,X)$, is to avoid recursive calls whenever the subtree rooted at $r$ cannot yield $\ell$-isolated maximal cliques.
First, we introduce a necessary condition for a clique grown from $C$ and $P$ to be $\ell$-isolated. Let $T$ be any maximal clique in $G[P]$ and let $t$ be its size. By the invariant property maintained by Algorithm˜1, $C\cup T$ is a clique in $G$. Such a clique has at least $\delta_{e}(C,P)+|C|(|P|-t)$ external edges: each of the |P|t|P|-t vertices not included in $C$ is indeed connected through an external edge with all the vertices in $C$. By definition, the number of external edges in an $\ell$-isolated clique must be strictly less than $\ell$ times the clique size. Hence, if
$\delta_{e}(C,P)+|C|(|P|-t)\geq\ell(|C|+t)$ (1)
then $C\cup T$ is not $\ell$-isolated. We now introduce a preliminary result that will be useful to prove the correctness of our pruning strategies.
Lemma 3.1.
Let $r$ be a node in $\mathcal{T}_{G}$ associated to a candidate set $C$ and a perspective set $P$. Let $P_{1}$ and $P_{2}$ be any two subsets of $P$ such that $|P_{1}|\leq|P_{2}|$. If $\delta_{e}(C,P)+|C|(|P|-|P_{2}|)\geq\ell(|C|+|P_{2}|)$, then both $C\cup P_{2}$ and $C\cup P_{1}$ are not $\ell$-isolated.
Proof 3.2.
As observed when we introduced Equation˜1, if $\delta_{e}(C,P)+|C|(|P|-|P_{2}|)\geq\ell(|C|+|P_{2}|)$, then $C\cup P_{2}$ is not $\ell$-isolated. The relationship $|P_{1}|\leq|P_{2}|$ yields the following chain of inequalities: $\delta_{e}(C,P)+|C|(|P|-|P_{1}|)\geq\delta_{e}(C,P)+|C|(|P|-|P_{2}|)\geq\ell(|C|+|P_{2}|)\geq\ell(|C|+|P_{1}|)$. Hence, $C\cup P_{1}$ is also not $\ell$-isolated.
Pruning by maximum cliques. The size of the largest clique in the subgraph $G[P]$ induced by $P$, i.e., its clique number $\omega(G[P])$, can be used to prune subtrees of $\mathcal{T}_{G}$. Namely, we prove in Lemma˜3.3 that it is correct to prune the subtree rooted at a node labelled $(C,P,X)$ whenever
$\delta_{e}(C,P)+|C||P|-\ell|C|\geq\omega(G[P])\cdot(\ell+|C|)$ (2)
Lemma 3.3.
Let $r$ be a node in $\mathcal{T}_{G}$ associated to a candidate set $C$ and a perspective set $P$. If $\delta_{e}(C,P)+|C||P|-\ell|C|\geq\omega(G[P])\cdot(\ell+|C|)$, then the subtree of $\mathcal{T}_{G}$ rooted at $r$ contains no leaf corresponding to an $\ell$-isolated maximal clique.
Proof 3.4.
We use Lemma˜3.1, choosing $P_{2}$ as a maximum clique in $G[P]$, so that $|P_{2}|=\omega(G[P])$. Notice that the inequality $\delta_{e}(C,P)+|C||P|-\ell|C|\geq\omega(G[P])\cdot(\ell+|C|)$ in the statement can be rewritten as $\delta_{e}(C,P)+|C|\big(|P|-\omega(G[P])\big)\geq\ell\big(|C|+\omega(G[P])\big)$. When it is satisfied, by Equation˜1 the subgraph induced by $C$ extended by a maximum clique of $G[P]$ cannot be $\ell$-isolated. By Lemma˜3.1, if we extend $C$ by any other clique of $G[P]$, the resulting subgraph is also not $\ell$-isolated. Any clique $P_{1}$ of $G[P]$ is indeed smaller than or equal to a maximum clique, i.e., $|P_{1}|\leq|P_{2}|=\omega(G[P])$ and thus Lemma˜3.1 can be applied.
Pruning by upper bounding the maximum clique size. Lemma˜3.3 implies that a check on $\omega(G[P])$ would be sufficient to prune at node $r$. Unfortunately, computing $\omega(G[P])$ is difficult — max-clique being $\mathsf{NP}$-hard. Hence, following an approach suggested in [DBLP:conf/kdd/WangCF13] for the computation of $\tau$-visible summaries, we use upper bounds to $\omega(G[P])$ to guide pruning. We will prove in Theorem˜3.5 that this maintains correctness with respect to our $\ell$-isolation problem. In particular, $\omega(G[P])$ cannot be larger than:
  1. 1.
    the number $|P|$ of vertices in $G[P]$;
  2. 2.
    the maximum degree of $G[P]$, denoted by $\Delta(G[P])$, augmented by $1$;
  3. 3.
    the maximum value of $k$ such that $G[P]$ contains at least $k$ vertices with degree $\geq k-1$, denoted by $\tau(G[P])$;
  4. 4.
    the degeneracy of $G[P]$, denoted by $\kappa(G[P])$, augmented by $1$. The degeneracy is the largest value of $k$ for which $G[P]$ has a $k$-core (i.e., a maximal subgraph in which each vertex has degree at least $k$).
The first upper bound $\omega(G[P])\leq|P|$ is the weakest, while the fourth one $\omega(G[P])\leq\kappa(G[P])$ is the strictest. In particular, the following chain of inequalities holds:
$\omega(G[P])\leq\kappa(G[P])+1\leq\tau(G[P])\leq\Delta(G[P])+1\leq|P|$ (3)
This can be intuitively explained as follows. The fact that $\omega(G[P])\leq\kappa(G[P])+1$ is well-known in the literature: in every clique of size $t$, every vertex has indeed degree t1t-1 and is thus part of a (t1)(t-1)-core, proving that the degeneracy is at least t1t-1. The value $\tau(G[P])$ represents a relaxed version of degeneracy, and thus $\kappa(G[P])+1\leq\tau(G[P])$. Indeed, the degeneracy $\kappa(G[P])$ can be computed by repeatedly removing a vertex of minimum degree from the graph, and updating the degree of remaining vertices along the way. Its relaxation $\tau(G[P])$ can be instead obtained by removing vertices based on their original degrees, without updates. Hence, it may be the case that more vertices satisfy the degree condition at point 3. Consider, for instance, a complete binary tree on four levels: $\tau(G[P])=4$, as the maximum degree 3 is attained by all the $6$ vertices in the two internal layers, while $\kappa(G[P])+1=2$, as the degeneracy of a tree is $1$. Finally, it can be easily seen that $\tau(G[P])\leq\Delta(G[P])+1$, because we cannot have $k$ vertices of degree larger than or equal to k1k-1 if $k\geq\Delta(G[P])+2$.
We use the upper bounds on $\omega(G[P])$ to obtain lower bounds on the total number of external edges that any maximal clique $T$, grown from the current clique $C$ and the candidate set $P$, can have. Namely, our algorithm prunes the computation at node $(C,P,X)$ whenever
$\delta_{e}(C,P)+|C||P|-\ell|C|\geq\overline{\omega}(G[P])\cdot(\ell+|C|)$ (4)
where $\overline{\omega}(G[P])$ can be any of the quantities involved in Equation˜3 (i.e., either $\omega(G[P])$ or any of its upper bounds). In the following we discuss the correctness of the pruning condition expressed by Equation˜4, proving that it is safe to prune using upper bounds $\overline{\omega}$ on $\omega$, instead of $\omega$ itself.
Theorem 3.5.
Given a graph $G$ and a positive integer $\ell$, let $\mathcal{T}_{G}$ be the search tree of Algorithm˜1 executed on $G$, let $r$ be a node in $\mathcal{T}_{G}$ associated to a candidate set $C$ and a perspective set $P$, and let $\overline{\omega}(G[P])$ be any upper bound on $\omega(G[P])$. If $\delta_{e}(C,P)+|C||P|-\ell|C|\geq\overline{\omega}(G[P])\cdot(\ell+|C|)$, then the subtree of $\mathcal{T}_{G}$ rooted at $r$ contains no leaf corresponding to an $\ell$-isolated maximal clique.
Proof 3.6.
We exploit Lemma˜3.1, choosing $P_{1}$ as a maximum clique in $G[P]$ and choosing $P_{2}$ as follows: if the upper bound $\overline{\omega}(G[P])$ in Equation˜3 is $|P|$, then $P_{2}$ is $P$ itself; if instead it is $\Delta(G[P])+1$, then $P_{2}$ is formed by any vertex of maximum degree in $P$ together with its neighbors; if it is $\tau(G[P])$, then $P_{2}$ is formed by any $\tau(G[P])$ vertices of $P$ each having degree at least $\tau(G[P])-1$; and when it is $\kappa(G[P])+1$, then $P_{2}$ is the $\kappa(G[P])$-core of $P$. Hence, $|P_{1}|=\omega(G[P])\leq\overline{\omega}(G[P])=|P_{2}|$. If the test in Equation˜4 is verified, by Lemma˜3.1 we have that the subgraph induced by $C$ extended by a maximum clique of $G[P]$ is not $\ell$-isolated. In particular, from the proof of Lemma˜3.1 we have $\delta_{e}(C,P)+|C|(|P|-\omega(G[P]))\geq\ell(|C|+\omega(G[P])$ (Equation˜2). This in turn implies, by Lemma˜3.3, that any other clique in $G[P]$ is not $\ell$-isolated and that pruning at $r$ does not entail discarding any $\ell$-isolated clique, thereby proving the statement.
We conclude with a remark on the running time: the quantities $|P|$, $\Delta(G[P])$, $\tau(G[P])$, and $\kappa(G[P])$ involved in Equation˜3 can be computed in $O(1)$, $O(1)$, $O(|P|)$, and $O(|E(G[P])|)$ time, respectively, yielding an inverse relation between running time and pruning power.
4 Experimental setup
Algorithms under evaluation. The four different heuristics discussed in Section˜3 are denoted in our analysis by size, degree, softcore, and degeneracy, respectively. We also consider a combined approach, called combo, that orderly integrates the size and softcore heuristics: we tested many different combinations of the four basic heuristics, and combo consistently proved to be the most effective one. For the sake of comparison, we include the algorithm proposed in [DBLP:journals/tcs/HuffnerKMN09], denoted by HKMN. We relied on the implementation used in [DBLP:journals/tcs/HuffnerKMN09]. Finally, we consider a baseline algorithm, derived from Algorithm˜1 [DBLP:conf/cocoon/TomitaTT04, DBLP:journals/tcs/TomitaTT06] and denoted by TTT, that filters each non-$\ell$-isolated maximal clique instead of reporting it.
Table 1: Statistics of the real-world benchmarks. $\Delta$ is the maximum degree, and $d$ is the degeneracy.
Graph ID Category Vertices Edges Max. Cliques $\Delta$ d
brightkite bk social 58,228 214,078 290,004 1,134 52
livemocha lm social 104,103 2,193,083 3,711,569 2,980 92
gowalla gw social 196,591 950,327 1,212,679 14,730 51
twitter ws social 465,017 833,540 805,170 677 30
youtube cy social 1,134,890 2,987,624 3,265,956 28,754 51
hyves hy social 1,402,673 2,777,419 2,590,217 31,883 39
ca-astroph ap co-authorship 18,771 198,050 36,427 504 56
dblp cd co-authorship 317,080 1,049,866 257,551 343 113
cfinder-google gc hyperlink 15,763 148,585 75,258 11,401 102
stanford sf hyperlink 281,903 1,992,636 1,055,936 38,625 71
italian-cnr ic hyperlink 325,557 2,738,969 1,425,378 18,236 83
notre-dame nd hyperlink 325,729 1,090,108 495,947 10,721 155
baidu bar hyperlink 415,624 2,374,044 4,596,143 127,066 228
web-google go hyperlink 875,713 4,322,051 1,417,580 6,332 44
cit-hepph phc citation 34,546 420,877 412,493 846 30
citeseer cs citation 384,054 1,736,145 1,232,951 1,739 15
Benchmarks. We conducted experiments on a variety of real-world and synthetic networks. Real-world networks, which are the main target of our analysis, are taken primarily from the KONECT repository (http://konect.cc) [DBLP:conf/www/Kunegis13]. Throughout this paper we focus on a representative subset of $16$ graphs, whose main properties are summarized in Table˜1. Our selection was guided by diversity criteria, taking into account graph size, number of maximal cliques, maximum degree $\Delta$, and degeneracy $d$. To further ensure heterogeneity, our benchmarks include different types of networks, namely social, hyperlink, co-authorship, and citation graphs. For the sake of completeness, the number of $\ell$-isolated maximal cliques for different values of $\ell$, as well as their percentage with respect to the total number of maximal cliques, is reported for these graphs in Table˜2.
Table 2: Evolution of the number of $\ell$-isolated maximal cliques and the percentage with respect to all maximal cliques for different values of $\ell$.
Graph $\ell=1$ $\ell=10$ $\ell=20$ $\ell=30$ $\ell=40$ $\ell=50$ $\ell=100$ $\ell=150$ $\ell=200$ $\ell=250$
brightkite 2346 32661 54118 67631 77577 85257 113328 181361 232791 264373
(0.81%) (11.26%) (18.66%) (23.32%) (26.75%) (29.40%) (39.08%) (62.54%) (80.27%) (91.16%)
livemocha 2 6619 32211 69514 112054 156226 361422 528104 668495 792579
(0.00%) (0.18%) (0.87%) (1.87%) (3.02%) (4.21%) (9.74%) (14.23%) (18.01%) (21.35%)
gowalla 3993 107537 176808 220693 252900 279548 376567 440791 511168 565420
(0.33%) (8.87%) (14.58%) (18.20%) (20.85%) (23.05%) (31.05%) (36.35%) (42.15%) (46.63%)
twitter 8 803 3401 6678 10471 15811 47327 82293 124235 603666
(0.00%) (0.10%) (0.42%) (0.83%) (1.30%) (1.96%) (5.88%) (10.22%) (15.43%) (74.97%)
youtube 41563 549757 756313 882061 975738 1049952 1291485 1450548 1566338 1662987
(1.27%) (16.83%) (23.16%) (27.01%) (29.88%) (32.15%) (39.54%) (44.41%) (47.96%) (50.92%)
hyves 44 829384 1218221 1295354 1394493 1515859 1920881 2018340 2113061 2190694
(0.00%) (32.02%) (47.03%) (50.01%) (53.84%) (58.52%) (74.16%) (77.92%) (81.58%) (84.58%)
ca-astroph 384 2629 5157 7462 9732 11941 21594 28811 32935 35077
(1.05%) (7.22%) (14.16%) (20.48%) (26.72%) (32.78%) (59.28%) (79.09%) (90.41%) (96.29%)
dblp 4762 103815 168752 205317 226139 237806 255468 257404 257537 257550
(1.85%) (40.31%) (65.52%) (79.72%) (87.80%) (92.33%) (99.19%) (99.94%) (99.99%) (100.00%)
cfinder-google 92 657 1097 1471 1597 1833 3679 5623 6858 17418
(0.12%) (0.87%) (1.46%) (1.95%) (2.12%) (2.44%) (4.89%) (7.47%) (9.11%) (23.14%)
stanford 1223 65655 107570 133984 153565 170323 217294 234531 249591 262519
(0.12%) (6.22%) (10.19%) (12.69%) (14.54%) (16.13%) (20.58%) (22.21%) (23.64%) (24.86%)
italian-cnr 1455 89276 141082 172788 193906 210687 267492 320474 369278 399686
(0.10%) (6.26%) (9.90%) (12.12%) (13.60%) (14.78%) (18.77%) (22.48%) (25.91%) (28.04%)
notre-dame 1292 108550 175070 211801 234167 251205 290741 317485 332918 345365
(0.26%) (21.89%) (35.30%) (42.71%) (47.22%) (50.65%) (58.62%) (64.02%) (67.13%) (69.64%)
baidu 22792 173799 306856 406504 466079 519940 631133 1031465 3280483 3301681
(0.50%) (3.78%) (6.68%) (8.84%) (10.14%) (11.31%) (13.73%) (22.44%) (71.37%) (71.84%)
web-google 8745 324503 535311 643214 709705 757338 888481 944254 975307 1003754
(0.62%) (22.89%) (37.76%) (45.37%) (50.06%) (53.42%) (62.68%) (66.61%) (68.80%) (70.81%)
cit-hepph 141 7273 23354 45525 71242 98142 251893 355771 390060 402261
(0.03%) (1.76%) (5.66%) (11.04%) (17.27%) (23.79%) (61.07%) (86.25%) (94.56%) (97.52%)
citeseer 10965 247124 478583 635481 746623 829906 1043603 1126876 1164634 1187575
(0.89%) (20.04%) (38.82%) (51.54%) (60.56%) (67.31%) (84.64%) (91.40%) (94.46%) (96.32%)
We also used synthetic instances generated according to two different models:
  • Random scale-free networks generated using a preferential attachment mechanism as proposed by Barabasi and Albert [BA02]. $\textrm{BA}_{n,m}$ networks are grown by sequentially adding new nodes, each with $m$ edges, that attach preferentially to existing nodes in proportion to their degrees. We used $n=100,000$ and $m\in\{25,50\}$.
  • Random networks based on the $G_{n,m,p}$ model (see [DBLP:journals/tcs/BehrischT06] and the references therein), which is also used in [DBLP:journals/tcs/HuffnerKMN09]. These networks have $n$ vertices; each of them is assigned $m$ features, each with probability $p$; and all the vertices with the same feature form a clique. On these instances, for the sake of reproducibility, we selected parameter values similar to [DBLP:journals/tcs/HuffnerKMN09], varying $n$ in $[50,550]$ with step $50$, $m$ in $[5,95]$ with step $10$, $p$ in $[0.025,0.25]$ with step $0.025$, and $\ell$ in $[10,200]$ with step $10$.
Performance metrics. In addition to running times, we profiled the number of recursive calls and contrasted these values with the number of $\ell$-isolated cliques. The running time of each heuristic depends indeed both on its pruning capabilities and on the overhead introduced by the pruning test in Equation˜4: the more complex the pruning strategy, the larger the test overhead.
Experimental platform. All the experiments were run on a laptop equipped with an Intel® Core™ i5-1135G7 CPU (4 cores, 4.2 GHz) and 8 GB of RAM. Our algorithms were implemented in OCaml (https://ocaml.org) to ensure a fair comparison with the available implementation used in [DBLP:journals/tcs/HuffnerKMN09].
5 Experimental results
In this section we present the key findings of our experimental analysis. Section˜5.1 examines the properties of $\ell$-isolated cliques in real-world benchmarks. We evaluate our pruning strategies in Section˜5.2, identifying the most effective ones, which are compared against state-of-the-art methods in Section˜5.3. We complement our study with results on synthetic instances in Section˜5.4.
5.1 On the distribution of $\ell$-isolated maximal cliques
Figure 1: $\ell$-isolated clique size distribution for different values of $\ell$.
In Figure˜1 we summarize the distribution of maximal cliques by size, together with their breakdown into $\ell$-isolated cliques for different values of $\ell$.
The upper profile of the stacked bars shows the total number of maximal cliques with a given number of vertices. In all benchmarks, most cliques consist of a few vertices: the number of cliques grows quickly for small sizes, then drops sharply. The maximum clique size varies substantially across graph types (cfr. Table˜1), reflecting different connectivity patterns: the first six social networks form only small cliques (typically less than 20 vertices), the two co-authorship networks and the two citation networks reach moderate sizes, while the six Web graphs contain the largest cliques, sometimes exceeding 150 vertices. The difference can be best appreciated on graphs of different types but similar size: compare, e.g., the maximum clique size in gowalla, dblp, and notre-dame, which have comparable numbers of edges.
This analysis can be refined by considering the color segments within each bar of Figure˜1, that show how many maximal cliques are $\ell$-isolated for several values of $\ell$. For small values of $\ell$ (e.g., $\ell\leq 20$), the isolation requirement is very strict and only a small fraction of small cliques qualify: large cliques are typically excluded, as they tend to be embedded in denser regions of the graph with many external connections. The distributions for small $\ell$ are thus sharply truncated and dominated by cliques with just a few vertices. As $\ell$ increases, the number of $\ell$-isolated cliques grows, being the property monotone (see Section˜2.1). In addition, larger values of $\ell$ allow increasingly larger cliques to appear in the summary, extending the tail of the distribution. For large thresholds, such as $\ell=200$, many more cliques are admitted: in many networks (e.g., brightkite, web-google, cit-hepph) the profile for 200-isolated cliques approaches that of the total distribution, making the isolation criterion almost ineffective.
5.2 Assessment of different pruning strategies
Figure 2: Comparison of the recursive calls (%) and the running time (% with respect to the slowest execution) of different pruning strategies for $\ell=50$.
We now assess the effectiveness of different pruning strategies considering two different metrics: number of recursive calls and running time. Because the results are largely consistent across different values of $\ell$, Figure˜2 shows only the case $\ell=50$. The left radar plot in Figure˜2 reports the percentage of recursive calls with respect to the number one would have without pruning. This gives a quantitative view of the pruning power of the heuristics: the larger the area, the weaker the pruning. In spite of a large variability across benchmarks, size is always the weakest, as expected from the theoretical analysis, producing the largest number of calls. On some benchmarks, e.g., twitter, this number can be much larger than the other strategies and very close to the baseline. As expected, combo and softcore overlap perfectly, since combo is size followed by softcore whenever size fails, thus inheriting the pruning power of softcore. While degeneracy consistently achieves the strongest pruning, yielding the smallest areas across all datasets, it is rather similar to softcore and combo.
Figure 3: Comparison of size, combo, HKMN, and TTT on real-world graphs.
Interestingly, the picture is reversed in the right radar plot, which reports the running time normalized with respect to the slowest execution: degeneracy, which is the most powerful pruning strategy, is also almost always the slowest, with the exception of baidu, where degree performs worst. The heuristics with weaker pruning, most notably size, softcore and combo, are instead the fastest. Among these, softcore and combo behave very similarly, with combo being slightly preferable in some cases. degree exhibits an intermediate performance across benchmarks. Overall, the results highlight a clear trade-off between pruning power and efficiency: the running times observed in theory to perform the pruning tests (see Section˜3) appear to have a stronger impact on performance than the actual amount of pruned nodes. Since running time is ultimately the main performance metric, in the remainder of this section we will focus on size and combo, which appear to be the most effective strategies.
5.3 Comparison with state-of-the-art competitors
In Figure˜3 we compare our best heuristics, size and combo, against HKMN and TTT, also analyzing how $\ell$ affects the running times. We chose a range of values for $\ell$ based on the observations in Section˜5.1, covering both the lower and the upper spectrum of the isolation factor. A few trends emerge clearly from the charts. First, the running time of TTT is independent of $\ell$, thus remaining constant. The two heuristics size and combo typically deliver the best performance over a large range of values of $\ell$, with runtimes that remain low and scale smoothly as $\ell$ increases, when larger portions of $\mathcal{T}_{G}$ need to be explored. Their curves are often close (see, e.g., livemocha or cfinder-google), combo is marginally faster than size on a few instances (most notably, stanford and italian-cnr), but size seems preferable in most of the other ones. On the other side, HKMN is by far the slowest algorithm when $\ell$ gets large, with runtimes that grow quickly with $\ell$: this is due to an $\ell$-parameterized subroutine for finding minimal vertex covers. It can be instead competitive when $\ell$ is very small. TTT is only occasionally competitive with size and combo, but still outperforms HKMN for large $\ell$ except on web-google and dblp. The observed relative differences between the algorithms persist across the datasets, showing that even weak pruning heuristics that add a small test overhead on nodes of $\mathcal{T}_{G}$ translate into significant speedups that can be as large as $4\times$ with respect to HKMN and $2\times$ with respect to TTT.
5.4 Experiments with synthetic networks
We now report experiments on synthetic instances that mostly confirm the analysis on real-world graphs.
(a) Comparison between HKMN and TTT.
(b) Comparison among size, combo, and TTT.
Figure 4: Results on BAn,m synthetic graphs. The dashed red line shows the number of $\ell$-isolated maximal cliques relative to the total number.
We first discuss experiments on BAn,m networks. Figure˜4 compares TTT against HKMN (Figure˜4(a)) and size and combo (Figure˜4(b)), respectively. For $m=25$ and $\ell\leq 20$, we have that HKMN is comparable with combo and both of them perform better than size. However, the dashed red line also shows that in the above range the $\ell$-isolated maximal cliques are very few (we also experimentally assessed their small size, akin to real-world networks). For higher values of $\ell$, the running time of HKMN seems to be strongly dependent on the output size. Heuristic size (and sometimes combo), instead, provide a more efficient solution when $\ell\geq 30$. Curve trends are similar when $m=50$. The percentage of $\ell$-isolated cliques decreases: for $\ell=250$ it is barely above $60\%$ and remains close to zero for $\ell\leq 50$ (in this range HKMN is the fastest algorithm). When $\ell>50$, TTT is faster than combo and HKMN, but size remains the preferred solution.
Networks generated according to the $G_{n,m,p}$ model have substantially different properties: they are typically very dense and $\ell$-isolated maximal cliques are barely present for small values of $\ell$ (e.g., $\ell<80$). Conversely, for high values of $\ell$ (e.g., $\ell>100$), their percentage becomes close to $100\%$, after a very steep transition (see the dashed red line in Figure˜5). We reproduced the analysis of [DBLP:journals/tcs/HuffnerKMN09], reporting in Figure˜5 only the results with respect to $\ell$. Our results confirm those in [DBLP:journals/tcs/HuffnerKMN09], with HKMN outperforming other approaches for small values of $\ell$. TTT seems instead the algorithm of choice for larger values. In any case, these graphs appear to represent worst-case scenarios for the algorithms proposed in this paper, although they remain very far from the properties of real-world networks.
Figure 5: Results on $G_{n,m,p}$ synthetic graphs. The dashed red line shows the number of $\ell$-isolated maximal cliques relative to the total number.
6 Conclusions and open problems
This paper addressed the problem of enumerating isolated cliques, introducing pruning heuristics that can be applied on top of classical maximal cliques enumeration algorithms. We have theoretically proved correctness and conducted an extensive experimental analysis on a variety of benchmarks, proving the effectiveness of two of our variants against state-of-the-art competitors. Overall, $\ell$-isolation appears to be a useful lens for filtering cliques according to how strongly they are separated from the rest of the network. Small values of $\ell$ expose only the most tightly isolated structures, at the cost of excluding larger cliques. At the other extreme, too large values of $\ell$ may contribute little beyond the raw clique distribution, as almost every clique qualifies. The most informative patterns emerge at intermediate thresholds, such as $\ell=50$ or $100$, which strike a balance between selectivity and inclusiveness: they admit a substantial fraction of cliques, even larger ones, while still preserving the notion of relative isolation. A natural way to address the tradeoff between selectivity and clique size, which is closely tied to the choice of $\ell$, could be to introduce alternative metrics in which the isolation factor is defined relative to clique size, rather than as an absolute threshold.
Isolation, and clique summarization more broadly, pose many other open problems (see also [DFP25]). In particular, it would be worthwhile to investigate whether our approach can be extended to the computation of $\ell$-isolated cliques by size (Problem 3 in [DFP25]) and to the stronger notion of max-$\ell$-isolation (see Section˜2.1). Another promising direction is to examine whether the problem admits fixed-parameter tractable algorithms under alternative parameters, such as degeneracy, either by refining the techniques of [DBLP:conf/esa/ItoIO05, DBLP:journals/talg/ItoI09] or by focusing on specific graph families with distinctive structural properties.

### v2

@@
 This follows immediately from the definition of $\ell$-isolation, since for an $\ell_{2}$-isolated set $I$ we have $\delta_{e}(I)<\ell_{2}\cdot k<\ell_{1}\cdot k$ when $\ell_{1}>\ell_{2}$.
-Hence, when searching for $\ell$-isolated maximal cliques, the larger is the value of $\ell$, the weaker is the isolation of the cliques relative to the rest of the graph, and the is larger the number of $\ell$-isolated cliques in the graph.
+Hence, when searching for $\ell$-isolated maximal cliques, the larger is the value of $\ell$, the weaker is the isolation of the cliques relative to the rest of the graph, and the larger is the number of $\ell$-isolated cliques in the graph.
 If $\ell$ is high enough, the set of $\ell$-isolated maximal cliques coincides with the set of all maximal cliques of $G$.
@@
 5:procedure ProcMCE($C,P,X$)
-6:  if $P=\varnothing$ and $X=\varnothing$ then return $C$
-7:  Choose a pivot vertex $v_{p}$ in $P\cup X$ with highest $|N(v_{p})\cap P|$
-8:  for all $v\in P\setminus N(v_{p})$ do
+6:  if $P=\varnothing$ and $X=\varnothing$ then
+7:   yield $C$
+8:   return
+9:  Choose a pivot vertex $v_{p}$ in $P\cup X$ with highest $|N(v_{p})\cap P|$
+10:  for all $v\in P\setminus N(v_{p})$ do
 […]
-10:   $P\leftarrow P\setminus\{v\}$
-11:   $X\leftarrow X\cup\{v\}$
+12:   $P\leftarrow P\setminus\{v\}$
+13:   $X\leftarrow X\cup\{v\}$
 When invoked on a graph $G$, Algorithm˜1 implicitly explores a search tree $\mathcal{T}_{G}$.
@@
 Otherwise, the maximality condition is not satisfied (vertices in $X$ could be added to $C$), the clique is discarded, but no recursive call needs to be done.
+Figure 1: A running example of Algorithm˜1 on a small graph.
+Green, blue, and red vertices belong to $P$, $C$, and $X$, respectively.
+The current pivot vertex $v_{p}$ is circled with a dotted curve.
+The neighbourhood of $v_{p}$ is circled with a dashed curve, the current vertex $v$ is circled with a red solid curve.
+A running example of Algorithm˜1 on a small graph is depicted in Figure˜1.
+When the first call of the procedure ProcMCE is launched, all the vertices are in $P$, while $X=C=\varnothing$ (Figure˜1(b)).
+The pivot vertex is $v_{p}=v_{2}$ (line 9 of Algorithm˜1) and the procedure iterates on vertex $v_{2}$ only, since all other vertices in $P$ are neighbors of $v_{p}$ (line 10 of Algorithm˜1).
+Hence, in the second call of ProcMCE (Figure˜1(c)) vertex $v_{2}$ has been moved to $C$, $v_{p}=v_{3}$, and line 10 of Algorithm˜1 only iterates on $v_{3}$.
+Instead, Call 3 iterates on both $v_{1}$ and $v_{4}$: the iteration on $v_{1}$ (Figure˜1(d)) launches Call 4 (Figure˜1(e)) while the iteration on $v_{4}$ (Figure˜1(f)) launches Call 5 (Figure˜1(g)).
+Call 4 and Call 5 detect each a maximal clique, since both $P$ and $X$ are empty (line 6 of Algorithm˜1).
+When Call 5 terminates, the control returns to Call 3 (Figure˜1(h)) and then to Call 2 (Figure˜1(i)) and Call 1 (Figure˜1(j)), terminating the computation.
 Algorithm˜1 requires $\mathcal{O}(3^{n/3})$ time [DBLP:journals/tcs/TomitaTT06].
@@
 Unfortunately, computing $\omega(G[P])$ is difficult — max-clique being $\mathsf{NP}$-hard.
-Hence, following an approach suggested in [DBLP:conf/kdd/WangCF13] for the computation of $\tau$-visible summaries, we use upper bounds to $\omega(G[P])$ to guide pruning.
+Hence, we exploit upper bounds to $\omega(G[P])$ to guide pruning, inspired by the approach suggested in [DBLP:conf/kdd/WangCF13] for the computation of $\tau$-visible summaries.
 We will prove in Theorem˜3.5 that this maintains correctness with respect to our $\ell$-isolation problem.
@@
 In addition to running times, we profiled the number of recursive calls and contrasted these values with the number of $\ell$-isolated cliques.
-The running time of each heuristic depends indeed both on its pruning capabilities and on the overhead introduced by the pruning test in Equation˜4: the more complex the pruning strategy, the larger the test overhead.
+The running time of each heuristic depends indeed both on its pruning capabilities and on the overhead introduced by the pruning test in Equation˜4: the more complex the pruning strategy is, the larger the test overhead.
 Experimental platform.
@@
 Figure 2: $\ell$-isolated clique size distribution for different values of $\ell$.
-In Figure˜1 we summarize the distribution of maximal cliques by size, together with their breakdown into $\ell$-isolated cliques for different values of $\ell$.
+In Figure˜2 we summarize the distribution of maximal cliques by size, together with their breakdown into $\ell$-isolated cliques for different values of $\ell$.
 The upper profile of the stacked bars shows the total number of maximal cliques with a given number of vertices.
@@
 The difference can be best appreciated on graphs of different types but similar size: compare, e.g., the maximum clique size in gowalla, dblp, and notre-dame, which have comparable numbers of edges.
-This analysis can be refined by considering the color segments within each bar of Figure˜1, that show how many maximal cliques are $\ell$-isolated for several values of $\ell$.
+This analysis can be refined by considering the color segments within each bar of Figure˜2, that show how many maximal cliques are $\ell$-isolated for several values of $\ell$.
 For small values of $\ell$ (e.g., $\ell\leq 20$), the isolation requirement is very strict and only a small fraction of small cliques qualify: large cliques are typically excluded, as they tend to be embedded in denser regions of the graph with many external connections.
@@
 We now assess the effectiveness of different pruning strategies considering two different metrics: number of recursive calls and running time.
-Because the results are largely consistent across different values of $\ell$, Figure˜2 shows only the case $\ell=50$.
-The left radar plot in Figure˜2 reports the percentage of recursive calls with respect to the number one would have without pruning.
+Because the results are largely consistent across different values of $\ell$, Figure˜3 shows only the case $\ell=50$.
+The left radar plot in Figure˜3 reports the percentage of recursive calls with respect to the number one would have without pruning.
 This gives a quantitative view of the pruning power of the heuristics: the larger the area, the weaker the pruning.
@@
 5.3 Comparison with state-of-the-art competitors
-In Figure˜3 we compare our best heuristics, size and combo, against HKMN and TTT, also analyzing how $\ell$ affects the running times.
+In Figure˜4 we compare our best heuristics, size and combo, against HKMN and TTT, also analyzing how $\ell$ affects the running times.
 We chose a range of values for $\ell$ based on the observations in Section˜5.1, covering both the lower and the upper spectrum of the isolation factor.
@@
 TTT is only occasionally competitive with size and combo, but still outperforms HKMN for large $\ell$ except on web-google and dblp.
-The observed relative differences between the algorithms persist across the datasets, showing that even weak pruning heuristics that add a small test overhead on nodes of $\mathcal{T}_{G}$ translate into significant speedups that can be as large as $4\times$ with respect to HKMN and $2\times$ with respect to TTT.
+The observed relative differences between the algorithms persist across the datasets.
+Even weak pruning heuristics that add a small test overhead on nodes of $\mathcal{T}_{G}$ translate into significant speedups, which can be as large as $4\times$ and $2\times$ with respect to HKMN and TTT, respectively.
 5.4 Experiments with synthetic networks
@@
 We first discuss experiments on BAn,m networks.
-Figure˜4 compares TTT against HKMN (Figure˜4(a)) and size and combo (Figure˜4(b)), respectively.
+Figure˜5 compares TTT against HKMN (Figure˜5(a)) and size and combo (Figure˜5(b)), respectively.
 For $m=25$ and $\ell\leq 20$, we have that HKMN is comparable with combo and both of them perform better than size.
@@
 Networks generated according to the $G_{n,m,p}$ model have substantially different properties: they are typically very dense and $\ell$-isolated maximal cliques are barely present for small values of $\ell$ (e.g., $\ell<80$).
-Conversely, for high values of $\ell$ (e.g., $\ell>100$), their percentage becomes close to $100\%$, after a very steep transition (see the dashed red line in Figure˜5).
-We reproduced the analysis of [DBLP:journals/tcs/HuffnerKMN09], reporting in Figure˜5 only the results with respect to $\ell$.
+Conversely, for high values of $\ell$ (e.g., $\ell>100$), their percentage becomes close to $100\%$, after a very steep transition (see the dashed red line in Figure˜6).
+We reproduced the analysis of [DBLP:journals/tcs/HuffnerKMN09], reporting in Figure˜6 only the results with respect to $\ell$.
 Our results confirm those in [DBLP:journals/tcs/HuffnerKMN09], with HKMN outperforming other approaches for small values of $\ell$.